News

Network status: New archive and pipeline; Spectral camera status; Sinistro redo

The 2015B semester is drawing to a close. The 2016A semester will begin on April 1. Results from LCOGT's time allocation process will be distributed soon.

1. Beginning with the start of the 2016A semester, LCOGT data products will be stored in a new in-house archive. The new archive has been under development throughout the 2015B semester, and we have now ingested all data back to May 1, 2014. The new archive - content and interfaces - are currently being tested. Please begin retrieving your data products from the new archive now.  Both raw and processed data are available.  The UI is available at https://archive.lcogt.net/ . The authentication system is the same as the one for the observatory portal. Documentation of the archive and the API are accessible from links at the top of the UI. If you encounter a problem or would like to make a suggestion, please click the "Feedback" link at the bottom of the UI. By the start of the new semester, we will redirect the website links to point to the new archive.  The IPAC archive will continue to be available until May 31, 2016, at which point it will be retired.  More information about the new archive will be available on the LCOGT website soon.

2. We are also in the process of testing and deploying a new data reduction pipeline. Since before the start of routine operations in 2014, we've been processing all frames with a customized version of Starlink's ORAC-DR pipeline. The new pipeline is being developed in-house, and its main advantage is that it will be easier to adapt and support than ORAC-DR. It should be faster, too. We expect to begin processing all incoming frames with the new pipeline by the start of the 2016A semester. Subsequently, we will reprocess all archived frames back to May 1, 2014.

3. The Spectral camera on Faulkes Telescope South (2m) has been out-of-service since February 22. The original problem was that a coolant line ruptured and leaked fluid onto the camera, causing the shutter to malfunction. The shutter was replaced on March 8, but now the camera window has condensation on it, and the autoguider camera is malfunctioning. The Spectral camera will remain offline until these problems are fixed. The FTS is still available for spectroscopic observations with FLOYDS.

4. The 1-meter telescope in Dome A at Siding Spring was fitted with a Sinistro camera on October 13. The 1-meter telescope in Dome C at Sutherland was fitted with a Sinistro camera on December 13. The commissioning of these new cameras has been arduous, and unfortunately, these cameras simply do not perform as well in the field as they did on our test telescope in Santa Barbara. The commissioning team has decided to replace these two cameras with the next two Sinistros that are available here in the lab.
Share on Twitter Share on Facebook